OpenStudy (anonymous):

i will medal The density of aluminum is 2.375 grams per cubic centimeter. Find the volume of a portion of a sheet of aluminum that has a mass of 0.78 kilograms? 3.04 cubic centimeters 1.8525 cubic centimeters 328.4 cubic centimeters 0.00304 cubic centimeters

OpenStudy (anonymous):

\[\frac{ Weight }{ Density } = Volume\] \[\frac{ 780g }{ 2.375\frac{ g }{ cm ^{3} } } = 328.421052632 cm ^{3}\]

OpenStudy (133):

Your answer is going to be 3.04 because when you do 2.375 divided by 0.78 you get 3.04 hope that helps.